OpenStudy (anonymous):

checking my answer: Solve y y' =1. Is y(x) = (2x+C)^(1/2) where C is arbitrary constants asolution y y' = 1! y(0) = 4

OpenStudy (anonymous):

\[y'=\frac{1}{\sqrt{2x+C}}\]

OpenStudy (anonymous):

so it looks like \(yy'=1\) no matter what

OpenStudy (anonymous):

y (dy/dx) = 1 y dy = dx (because y(0) = 4) int y dy (4 from y) = x dx (0 from x) (y^2/2) - (4^2/2) = x ((y^2)/2) - 8 = x y^2 = (x+8)2 y^2 = 2x + 16

OpenStudy (anonymous):

checking (y(0) = 4) y^2 = 2x + 16 y^2 = 2(0) + 16 y^2 = 16 y = sqrt(16) y = 4